OpenStudy (anonymous):

Which is a factor of 15xy – 45x – 6y + 18

OpenStudy (anonymous):

15xy – 45x – 6y + 18 = 3 (5xy - 15x - 2y + 6) = 3 ( 5x(y - 3) - 2 (y - 3) ) = 3 (5x-2) (y - 3). This is known as Simon's Favorite Factoring Trick (SFFT)!